# What You See Is What Matters: A Novel Visual and Physics-Based Metric for Evaluating Video Generation Quality

### v1

@@ Abstract
As video generation models advance rapidly, assessing the quality of generated videos has become increasingly critical. Existing metrics, such as Fréchet Video Distance (FVD), Inception Score (IS), and ClipSim, measure quality primarily in latent space rather than from a human visual perspective, often overlooking key aspects like appearance and motion consistency to physical laws. In this paper, we propose a novel metric, VAMP (Visual Appearance and Motion Plausibility), that evaluates both the visual appearance and physical plausibility of generated videos. VAMP is composed of two main components: an appearance score, which assesses color, shape, and texture consistency across frames, and a motion score, which evaluates the realism of object movements. We validate VAMP through two experiments: corrupted video evaluation and generated video evaluation. In the corrupted video evaluation, we introduce various types of corruptions into real videos and measure the correlation between corruption severity and VAMP scores. In the generated video evaluation, we use state-of-the-art models to generate videos from carefully designed prompts and compare VAMP’s performance to human evaluators’ rankings. Our results demonstrate that VAMP effectively captures both visual fidelity and temporal consistency, offering a more comprehensive evaluation of video quality than traditional methods.
@@ 2 Related Work
Image Generation The field of image generation has seen remarkable progress in recent years, primarily driven by advances in deep learning techniques. Generative Adversarial Networks (GANs) [8] have been at the forefront of this revolution, with architectures like StyleGAN [12] and BigGAN [1] pushing the boundaries of image quality and diversity. More recently, diffusion models [10] have emerged as a powerful alternative, demonstrating exceptional image quality and flexibility in tasks such as text-to-image generation [20][4]. These advancements in image generation have laid the groundwork for tackling the more complex task of video generation, providing crucial insights into high-dimensional data synthesis and the importance of perceptual quality in generated content.
@@
Video Generation Building upon the success of image generation models, video generation has become an active area of research, presenting unique challenges due to the temporal dimension. Early approaches extended GANs to the video domain, with models like VGAN [25] and MoCoGAN [23] attempting to capture both motion and content. Recent work has explored the use of 3D convolutions and attention mechanisms, as seen in Video Transformer Network [16], to better model long-range dependencies in video sequences. The advent of large language models has also influenced video generation, with text-to-video models like Make-A-Video [15] demonstrating impressive results in generating videos from textual descriptions. Exploration on using diffusion models led to improvement on text-guided video generation quality. [33]. Despite these advancements, consistent long-term motion and adherence to physical laws remain significant challenges in the field of video generation.
@@
Metrics for Video Generation As video generation models have evolved, so too has the need for robust evaluation metrics. Traditional image quality metrics like Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) have been adapted for video [29], but they often fail to capture the complexities of human perception in video quality assessment. The Fréchet Video Distance (FVD) [24], an extension of the Fréchet Inception Distance (FID) used in image generation, has become a popular metric for evaluating the quality and diversity of generated videos. More recently, perceptual metrics leveraging large pre-trained models, such as CLIP score [19], have been applied to video evaluation, offering a more semantic assessment of generated content. Leveraging multi-modal language model and large-scale human feedback to develop evaler demonstrate another promising direction. [9] However, these metrics still struggle to fully capture aspects such as temporal consistency, physical plausibility, and long-term coherence in generated videos. Our work aims to address these limitations by introducing a novel metric that specifically targets these crucial aspects of video quality.
@@ 3.1 Point Sampling and Mask Selection
To properly track and identify the region of interest in the image to evaluate, we use the above mentioned point sampling methods to identify the important regions to track. We then use SAM2 to track the movement of those important regions. We mainly experimented with two selection methods: 1) SAM2-based sampling and 2) SIFT-based Sampling, which have trade-off between quality and efficiency.
@@ 3.2.1 Appearance Score
Color Similarity. Color similarity is evaluated by comparing the color distribution between consecutive frames. To quantify this, we employ the Earth Mover’s Distance (EMD), which is also known as Wasserstein distance, to measure the dissimilarity between the color distributions between corresponding objects:
@@ 4.2 Generated Video Evaluation
Experiment Setup. For this evaluation, we sampled 3,000 prompts from the VidProM dataset, which includes a diverse range of video prompts and scenarios. Four video generation models were used to produce these videos: Pika [17], VideoCraft2 [3] , Text2Video-Zero [14] , and ModelScope [26] . The sampled prompts were used by the video generation models to generate corresponding videos. With the generated videos, we evaluate them using our proposed metric, calculating both the appearance score and motion score for each video. The evaluation provides insights into how well the generated videos maintain visual fidelity and motion coherence across different models and scenarios.
@@ 5 Results and Analysis
Figure 4: Normalized changes in metric values across corruption types and levels relative to the original video. Each heatmap represents the impact of a specific corruption type on four evaluation metrics (ClipSim, IS, FVD, and VAMP). Blue indicates a positive correlation, while red indicates a negative correlation.

### v2

@@ Abstract
As video generation models advance rapidly, assessing the quality of generated videos has become increasingly critical. Existing metrics, such as Fréchet Video Distance (FVD), Inception Score (IS), and ClipSim, measure quality primarily in latent space rather than from a human visual perspective, often overlooking key aspects like appearance and motion consistency to physical laws. In this paper, we propose a novel metric, VAMP (Visual Appearance and Motion Plausibility), that evaluates both the visual appearance and physical plausibility of generated videos. VAMP is composed of two main components: an appearance score, which assesses color, shape, and texture consistency across frames, and a motion score, which evaluates the realism of object movements. We validate VAMP through two experiments: corrupted video evaluation and generated video evaluation. In the corrupted video evaluation, we introduce various types of corruptions into real videos and measure the correlation between corruption severity and VAMP scores. In the generated video evaluation, we use state-of-the-art models to generate videos from carefully designed prompts and use VAMP to compare the models’ performances. Our results demonstrate that VAMP effectively captures both visual fidelity and temporal consistency, offering a more comprehensive evaluation of video quality than traditional methods.
@@ 2 Related Work
Image Generation The field of image generation has seen remarkable progress in recent years, primarily driven by advances in deep learning techniques. Generative Adversarial Networks (GANs) [13] have been at the forefront of this revolution, offering a wide range of applications [14, 32, 21, 19, 53, 48, 5, 8]. Architectures such as StyleGAN [22] and BigGAN [3] have pushed the boundaries of image quality and diversity. More recently, diffusion models [16, 55, 36, 1, 56] have emerged as a powerful alternative, demonstrating exceptional image quality and flexibility in tasks such as text-to-image generation [34][9]. These advancements in image generation have laid the groundwork for tackling the more complex task of video generation, providing crucial insights into high-dimensional data synthesis and the importance of perceptual quality in generated content.
@@
Video Generation Building upon the success of image generation models, video generation has become an active area of research, presenting unique challenges due to the temporal dimension. Early approaches extended GANs to the video domain, with models like VGAN [42] and MoCoGAN [40] attempting to capture both motion and content. Recent work has explored the use of 3D convolutions and attention mechanisms, as seen in Video Transformer Network [26], to better model long-range dependencies in video sequences. The advent of large language models has also influenced video generation, with text-to-video models like Make-A-Video [25] demonstrating impressive results in generating videos from textual descriptions. Exploration on using diffusion models led to improvement on text-guided video generation quality [57, 18, 17, 52, 2, 29, 50]. Despite these advancements, consistent long-term motion and adherence to physical laws remain significant challenges in the field of video generation.
@@
Metrics for Video Generation As video generation models have evolved, so too has the need for robust evaluation metrics. Traditional image quality metrics like Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) have been adapted for video [47], but they often fail to capture the complexities of human perception in video quality assessment. The Fréchet Video Distance (FVD) [41], an extension of the Fréchet Inception Distance (FID) used in image generation, has become a popular metric for evaluating the quality and diversity of generated videos. More recently, perceptual metrics leveraging large pre-trained models, such as CLIP score [33], have been applied to video evaluation, offering a more semantic assessment of generated content. Leveraging multi-modal language model and large-scale human feedback to develop evaler demonstrate another promising direction [15]. Some recent emergent metrics take a similar approach as FVD, measuring distances in the latent space with a more detailed focus on specific classes of characteristics [20, 28, 7, 27]. However, they still struggle to fully capture aspects such as temporal consistency, physical plausibility, and long-term coherence in generated videos. Our work aims to address these limitations by introducing a novel metric that specifically targets these crucial aspects of video quality.
@@ 3.1 Point Sampling and Mask Selection
To properly track and identify the region of interest in the image to evaluate, we use the above mentioned point sampling methods to identify the important regions to track. We then use SAM2 [35] to track the movement of those important regions. We mainly experimented with two selection methods: 1) SAM2-based sampling and 2) SIFT-based Sampling, which have trade-off between quality and efficiency.
@@ 3.2.1 Appearance Score
Color Similarity. Color similarity is evaluated by comparing the color distribution between consecutive frames. To quantify this, we employ the Earth Mover’s Distance (EMD) [37], which is also known as Wasserstein distance, to measure the dissimilarity between the color distributions between corresponding objects:
@@ 4.2 Generated Video Evaluation
Experiment Setup. For this evaluation, we sampled 3,000 prompts from the VidProM dataset [44], which includes a diverse range of video prompts and scenarios. Four video generation models were used to produce these videos: Pika [30], VideoCraft2 [6] , Text2Video-Zero [24] , and ModelScope [43] . The sampled prompts were used by the video generation models to generate corresponding videos. With the generated videos, we evaluate them using our proposed metric, calculating both the appearance score and motion score for each video. The evaluation provides insights into how well the generated videos maintain visual fidelity and motion coherence across different models and scenarios.
@@ 5 Results and Analysis
Figure 4: Normalized changes in metric values across corruption types and levels. Each heatmap represents the impact of a specific corruption type on four evaluation metrics (ClipSim, IS, FVD, and VAMP). Blue indicates a positive correlation, while red indicates a negative correlation.
@@ 6 Conclusion
Supplementary Material
Appendix A Experiment Details
We utilized the UCF-101 dataset to generate corrupted datasets for evaluation. For benchmarking against state-of-the-art generative models, we randomly sampled 3,000 videos from VidProM [44]. To ensure consistency, the prompts used for generating the videos were identical across all models.
During score calculation, there were instances where SAM2 failed to produce any segmentation masks. In such cases, we assigned a score of 0 to the corresponding videos. This decision was based on our observations that segmentation failures typically occur for videos of very low quality, particularly those with poor resolution.
For SIFT-based experiments, the weight tuple (0.3, 0.05, 0.05, 0.6) was applied for the final score calculation, corresponding to the components (color, shape, texture, motion). For SAM2-based experiments, we employed a weight tuple of (0.069, 0.138, 0.092, 0.7), ensuring a balanced contribution of appearance and motion components.
In terms of hardware, we primarily utilized NVIDIA A100 GPUs as accelerators. The inference tests for SAM2 were conducted on a single A100 GPU, using the original SAM2-large (sam2-hiera-large) implementation in a single-process setup. Conversely, SIFT inference tests were performed on a machine equipped with two CPU cores, demonstrating the computational efficiency of the SIFT-based approach.
Appendix B Point Sampling Algorithms
Method Brightness Gaussian Noise
0 1 2 3 4 5 0 1 2 3 4 5
VAMP-A $\uparrow$ 0.912 0.910 0.914 0.909 0.908 0.909 0.912 0.798 0.807 0.797 0.793 0.786
VAMP-M $\uparrow$ 0.578 0.569 0.571 0.564 0.557 0.559 0.578 0.508 0.518 0.526 0.507 0.499
VAMP $\uparrow$ 0.712 0.706 0.708 0.702 0.697 0.699 0.712 0.624 0.634 0.634 0.621 0.614
Impulse Noise Defocus Blur Black Shapes
0 1 2 3 4 5 0 1 2 3 4 5 0 1 2 3 4 5
0.912 0.918 0.918 0.756 0.891 0.887 0.912 0.918 0.919 0.901 0.899 0.882 0.912 0.900 0.888 0.884 0.869 0.861
0.578 0.545 0.544 0.432 0.421 0.425 0.578 0.601 0.603 0.570 0.537 0.507 0.578 0.550 0.499 0.466 0.468 0.471
0.712 0.694 0.694 0.562 0.609 0.610 0.712 0.728 0.729 0.703 0.682 0.657 0.712 0.690 0.655 0.634 0.628 0.627
Table 3: VAMP score results using the SIFT sampling method across different corruption types (Brightness, Gaussian Noise, Impulse Noise, Defocus Blur, and Black Shapes) at levels 0–5. The table presents scores for three components: VAMP-A (appearance-only), VAMP-M (motion-only), and the combined VAMP score.
Figure 5: Heatmap of Percentage Changes in VAMP Scores from Level 0 Across Corruption Types Using SIFT Sampling. This figure visualizes the percentage change in VAMP-A (appearance-only), VAMP-M (motion-only), and VAMP (combined) scores across five corruption types: Brightness, Gaussian Noise, Impulse Noise, Defocus Blur, and Black Shapes. Each subplot represents the changes in scores at corruption levels L1 to L5 relative to the baseline at L0.
Effective point sampling is essential for generating relevant masks, which are used for tracking and segmentation of objects in the VAMP pipeline. This section details the four implemented sampling algorithms: Random Sampling, Uniform Grid Sampling, SIFT Sampling and SAM Sampling.
B.1 Random Sampling
Random sampling generates a set of points by uniformly sampling coordinates within the dimensions of the input frame. This method is straightforward and computationally efficient, requiring no prior knowledge or feature extraction. The algorithm ensures diversity by randomly selecting $x$- and $y$-coordinates, effectively covering various regions of the frame.
Although simple, random sampling may lack contextual relevance, as it does not consider the visual content or structure of the frame. However, it is particularly useful as a baseline or fallback method when other algorithms, such as SIFT or SAM, fail to produce reliable points due to insufficient features or challenging input conditions. This method can also be combined with visualization to ensure that the sampled points are uniformly distributed across the frame.
B.2 Uniform Grid Sampling
Uniform grid sampling divides the frame into a grid-like structure and generates points at the center of each grid cell. This method ensures consistent spatial coverage, making it ideal for tasks that require uniform distribution of points across the frame. The number of grid cells in the $x$- and $y$-directions is determined by the specified number of points ($n_{x}$ and $n_{y}$), which can be adjusted to balance precision and computational cost.
Uniform grid sampling is particularly suited for scenarios where uniform coverage is essential, such as evaluating global properties of an image or video frame. Unlike random sampling, this method ensures predictable and repeatable point placement. However, it does not adapt to the content of the frame and may miss salient features if they are not aligned with the grid. Its simplicity and consistency make it a reliable choice for structured sampling tasks.
B.3 SIFT Sampling
The SIFT sampling method utilizes handcrafted features to identify key points within the input frame. The algorithm detects features by converting the input frame to grayscale and applying the SIFT detector. The resulting key points are returned as sampled points, representing visually salient features based on local gradients and intensity patterns.
For applications requiring clustered feature points, the SIFT-based sampling can be extended with DBSCAN (Density-Based Spatial Clustering of Applications with Noise). The SIFT Cluster algorithm groups detected key points into clusters and extracts their centroids, providing a reduced yet meaningful subset of feature points. This ensures that the sampled points represent the spatial distribution of features in the frame effectively.
The SIFT sampling methods are particularly suited for scenarios where computational efficiency is prioritized, as they rely on traditional feature extraction techniques without the need for pre-trained models.
B.4 SAM Sampling
The SAM-based sampling algorithm leverages the Segment Anything Model (SAM) architecture for point generation. The method utilizes the SAM2AutomaticMaskGenerator, a mask-generation module built on top of the SAM2 framework. This approach identifies meaningful regions in the input frame, extracting key points from the mask coordinates. Each point represents a salient feature, ensuring the sampled points are contextually significant.
This method operates by generating masks for the input frame using the SAM2 model, followed by extracting representative coordinates for each mask. The algorithm supports both CPU and GPU execution, with optimizations such as ‘torch.autocast‘ to improve efficiency. SAM sampling excels in scenarios where precise feature localization is critical, as it relies on learned priors from the SAM model to identify features of interest.
B.5 Efficiency and Performance Tradeoff between SIFT sampling and SAM sampling
Table 1 compares the results obtained using the SAM sampling method, while Table 3 presents the VAMP scores derived from the SIFT sampling method. A key observation is the tradeoff between efficiency and performance between these two approaches.
Figure 6: Sensitivity of VAMP Scores to Appearance Weights ($\alpha$) Across Corruption Types. Each subplot corresponds to a specific corruption type (Brightness, Gaussian Noise, Impulse Noise, Defocus Blur, and Black Shapes). The x-axis represents the corruption level (0–5), while the y-axis indicates the value of $\alpha$, the weight of the appearance score in the VAMP metric. The heatmap illustrates how varying $\alpha$ impacts the VAMP score at different levels of corruption for each type.
The SIFT sampling method demonstrates a significant advantage in computational efficiency, with an average sampling time of only 0.095 seconds per frame, compared to 2.69 seconds per frame for SAM2 sampling. This substantial difference makes SIFT sampling a more viable option for applications where rapid inference is required.
Table 4: Comparison of Sampling Efficiency Between SIFT and SAM2 Methods
Method Time (s/fr) Efficiency
SIFT 0.095 High
SAM2 2.69 Low
However, this efficiency comes at the cost of performance. The masks generated using SIFT sampling are less representative of the video content compared to those produced by SAM2, leading to less reliable VAMP scores. Specifically, as shown in Table 3, the correlations between corruption levels and VAMP scores are not as ideal, indicating that the SIFT sampling method struggles to capture the impact of corruption types effectively. In contrast, Table 1 demonstrates that the SAM2 sampling method yields VAMP scores with stronger and more consistent correlations to corruption levels, reflecting its superior ability to represent the overall video quality.
In summary, while SIFT sampling offers notable speed advantages, it sacrifices the representativeness and reliability of the resulting VAMP scores. For tasks that prioritize accurate and meaningful evaluations of video quality, SAM2 sampling is the preferred method despite its higher computational cost. Future work could explore hybrid approaches to balance efficiency and performance.
Appendix C VAMP Weight Sensitivity Analysis
Weights play a critical role in determining the VAMP score, as they balance the contributions of various components to the overall evaluation. Adjusting these weights emphasizes different characteristics of video quality, such as color, shape, texture, motion, or overall coherence. To better understand the impact of these weights, we present an analysis of VAMP’s weight sensitivity. We analyze the sensitivity of the overall VAMP score by varying the weights between the appearance score and motion score.
C.1 Sensitivity of VAMP Component Weights
The VAMP score ($S_{\text{VAMP}}$) combines appearance ($S_{\text{appearance}}$) and motion ($S_{\text{motion}}$) scores, weighted by $\alpha$ and $\beta$, to evaluate video quality. These weights are pivotal to the VAMP metric as they balance the contributions of appearance and motion components, allowing task-specific prioritization and making VAMP adaptable to diverse evaluation scenarios. However, this flexibility necessitates careful weight calibration to ensure meaningful and representative evaluations. To understand the sensitivity of VAMP scores to these weights, we analyze their impact across various corruption types and generative models. Figure 6 examines the effect of varying the appearance weight ($\alpha$) on VAMP scores under five corruption types: Brightness, Gaussian Noise, Impulse Noise, Defocus Blur, and Black Shapes. Figure 7 explores how $\alpha$ influences the VAMP scores of three generative models (VC2, MS, and Pika). Below, we discuss the findings from each figure, followed by a summary of key insights.
Figure 6 illustrates the sensitivity of VAMP scores to $\alpha$ across different corruption types. For Brightness, VAMP scores remain relatively stable across all corruption levels, suggesting that appearance weight variations have minimal impact. In contrast, Gaussian Noise and Impulse Noise exhibit sharper declines in VAMP scores as corruption levels increase, especially when $\alpha$ is low, highlighting the importance of motion consistency in these scenarios. Defocus Blur shows a significant drop in VAMP scores at higher corruption levels, particularly when $\alpha$ is large, indicating that motion plays a crucial role in mitigating its effects. Black Shapes exhibit substantial score reductions at higher corruption levels regardless of $\alpha$, underscoring the intrinsic difficulty of this corruption type.
Figure 7: Impact of $\alpha$ on VAMP Scores Across Different Video Generation Models. The heatmap compares the VAMP scores for different generative models (VC2, MS, and Pika) as a function of $\alpha$, the weight of the appearance score.
Figure 7 analyzes the impact of $\alpha$ on the VAMP scores of three generative models, revealing varying sensitivities to appearance and motion components. VC2 demonstrates consistent score improvements as $\alpha$ increases, indicating a stronger dependence on appearance components. MS (ModelScope) shows moderate sensitivity to $\alpha$, with a slower increase in scores, reflecting a more balanced reliance on both appearance and motion. Pika consistently outperforms the other models across all $\alpha$ values, showcasing its robustness in integrating both appearance and motion. Notably, Pika has the smallest sensitivity to $\alpha$ (0.032 difference), while MS exhibits the largest sensitivity (0.064 difference). These differences indicate that different models prioritize appearance and motion scores differently, with VC2 and MS achieving higher appearance scores and Pika excelling in motion consistency. A more even VAMP score across $\alpha$ reflects small deviations between appearance and motion scores, while a higher total VAMP value indicates superior overall quality. Pika’s high scores and minimal sensitivity suggest it achieves a well-balanced and high-quality output, making it a strong candidate for robust video generation tasks.
In summary, the sensitivity analysis reveals that VAMP scores are highly dependent on the corruption type and the generative model. Corruptions such as Gaussian Noise and Impulse Noise benefit from higher motion weighting (lower $\alpha$), while higher appearance weighting ($\alpha>0.5$) is essential for addressing Brightness and Black Shapes. Among the models, Pika demonstrates strong robustness across a wide range of $\alpha$, making it well-suited for diverse video generation tasks. Future work could focus on adaptive mechanisms to dynamically tune $\alpha$ based on the corruption type or specific evaluation requirements.